OpenStudy (anonymous):

for the pair of functions, f(x)=x^2+2 and g(x)=x+5, find the following a).(f*g)(x) b).(f*g)(2) c).(g*f)(x) d).(g*f)(2)

OpenStudy (id21):

f*g= \[f*g(x)=(x ^{2}+2)(x+5)=x^{3}+5x^{2}+2x+10\]

OpenStudy (id21):

\[(f \times g)(2); x=2; 2^{3}+5 \times 2^{2}+ 2\times2+10 = 8+20+4+10=42\]

OpenStudy (id21):

\[(g \times f)(x)=(x+5)(x^{2}+2)= x^{3}+2x+5x^{2}+10\]

OpenStudy (id21):

\[(g \times f) = 42\]

OpenStudy (id21):

Maybe I don't get something, but \[a \times b \] is \[b \times a\]

OpenStudy (anonymous):

thank u